# Central Answer Modeling for an Embodied Multi-LLM System

### v4

@@ Introduction
  • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds. To initial the debate we provide each LLM-based agent context with the following system prompt.
@@
    When prompted for the final answer at the end of the debate. The user prompt is Please give your final ”Yes/No” answer. Give a definite answer.
We then collect the final answers from each agent and take the majority vote as the final prediction.
Language-Guided Exploration
Figure 3: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment on a Matterport environment with 215 nodes. We report the mean and standard deviation over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
@@
Table 2: Performance measured in accuracy (Eq 1) of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment in $8$ different environments over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy. We repeated the experiment in $4$ of the environments but turned one of the agent malicious by providing responses opposite to what it believes are the correct answers. In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
We have all the LLM-based agents explore the household environment for a set amount of time before answering questions. In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala, Mullen Jr, and Manocha (2023). In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
@@
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in the Approach Section with the Question-Answering and Central Model Training phases. We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds). Experiments are ran on an Apple M1 Pro and macOS 14.5.
Results: The performance of the various CAM methods and baselines in the $8$ graph environments is shown in Table 2. We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities. From the table, we see CAM model consistently outperforms both baselines in accuracy (Eq 1. We report the average accuracy over $5$ seeds. We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers. Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method can best fit the generated query-answer data.
@@
where $i$ is index of the query parameterized by $(o,r)$. Table 3 reports the agreement over $5$ trials and we see that XGBoost has learned not to agree with the agents as much as MV and debate to attain its high accuracy. This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent. We further analyze the independent answers impact on training in the Feature Importance Analysis Section.
@@
Malicious Agent: We also show the benefits of CAM where one of the three agents is malicious. To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa. We then train all CAM methods and execute baselines regularly. Thus, the malicious agent contributes an answer it believes to be incorrect. From Table 2, we see that our CAM methods still outperform the MV and debate baselines. We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct. In the next section, we provide a feature importance of CAM.
Feature Importance Analysis
In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer. For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance (Breiman 2001; Fisher, Rudin, and Dominici 2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
@@
Conclusion and Limitations
In this paper, we tackle a variation of Embodied Question Answering with a Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer. While other methods have used a central module to aggregate independent agent answers for QA tasks, we specifically investigate this method in the context of embodied exploration in a household environment. We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query. We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating. A key benefit of our system is that it requires no communication during inference, reducing time costs. Furthermore, our approach can be combined with SOTA methods for exploration in unknown households. We also show that CAM is more robust to malicious agents than the majority vote and debate baslines. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions.
Appendix A Appendix
Appendix B Code and Dataset
@@ Appendix C Multi-LLM vs LLM
Appendix D Additional Visualization of Results
@@
Appendix E Decision Tree Visualization
In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section Feature Importance Analysis. This tree in Figure 11 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents. One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations. We see that the main feature used was actually the object query. We provide the object encoding labels in Table 6.
Figure 11: Visualization of Decision Tree from Section Feature Importance Analysis.

### v5

@@ 3.4.1 Non-learning-based Aggregation
  • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds. To initialize the debate we provide each LLM-based agent context with the following system prompt.
@@
    When prompted for the final answer at the end of the debate. The user prompt is Please give your final "Yes/No" answer. Give a definite answer. We then collect the final answers from each agent and take the majority vote as the final prediction.
3.4.2 Central Answer Model (CAM)
To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No" output. The model input dimension is $2+K$, where $K$ is the number of agents. The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$. The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$. $f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$. During inference, we input unseen questions and agent answers into the network to output a test-time answer. Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
We produce CAM models with various machine-learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), and Logistic Regression (LR). The NN was implemented with Pytorch using 3 linear layers. The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1. XGBoost used the default settings from the xgboost library package. RF from the sklearn library used $1000$ trees, and all other methods used the default sklearn settings.
4 Experimentation
We now describe our experimentation in detail. With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section). In all experiments we use GPT-4-turbo.
4.1 Query Generation
We extract the embodied question bank from the Matterport3D environments. For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm Li et al. (2022). The items space observed by GLIP can be found in the Appendix in Table 5.
For each item and room combination, we formulate the question using the template described in Section 3.3. Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes".
To create questions with ground-truth answers “No", for each query generated with a “Yes" ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
Note: Our setup is highly dependent on the object detector used. Error can thus be introduced by the object detection with GLIP and then the LLM answer generation. Some outputs of GLIP included toilets in the living room and a TV in the hallway. When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent. More error analysis on agent answers will be studied further in the Section 5.
4.2 Evaluation
We use a SOTA algorithm, LGX Dorbala et al. (2023) for the exploration phase.
Metric. To measure and compare the performances of our various MELE aggregation methods, we calculate the inference time accuracy, or correctness, of the model:
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from the given aggregation method.
4.3 Language-Guided Exploration
@@
Table 1: Performance measured in accuracy (Eq 1) of MELE with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment in $8$ different environments over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy. We repeated the experiment in $4$ of the environments but turned one of the agent malicious by providing responses opposite to what it believes are the correct answers. In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
Figure 2: Performance of MELE methods with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment on a Matterport environment with 215 nodes. We report the mean and standard deviation over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
We have all the LLM-based agents explore the household environment for a set amount of time before answering questions. In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al. (2023). In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
@@
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in Section 3 with the Question-Answering and Aggregation phases. We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds). Experiments were run on an Apple M1 Pro and macOS 14.5.
Results: The performance of the various CAM methods and non-learning aggregation methods in the $8$ graph environments is shown in Table 1. We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities. From the table, we see CAM model consistently outperforms both MV and debating in accuracy (Eq 1. We report the average accuracy over $5$ seeds. We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers. Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method best fits the generated query-answer data.
@@
where $i$ is index of the query parameterized by $(o,r)$. Table 2 reports the agreement over $5$ trials and we see that XGBoost has learned not to agree with the agents as much as MV and debate to attain its high accuracy. This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent. We further analyze the independent answers’ impact on training in the Section 5.
@@
Malicious Agent: We also show the benefits of CAM where one of the three agents is malicious. To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa. We then train all CAM methods and execute MV and debating regularly. Thus, the malicious agent contributes an answer it believes to be incorrect. From Table 1, we see that our CAM methods still outperform MV and debate. We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct.
5 Feature Importance Analysis
In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer. For our MELE system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of the CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al. (2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
@@ 5 Feature Importance Analysis
6 Conclusion
In this paper, we use Multi-Embodied LLM Explorers (MELE) to understand the effects of embodied exploration of multiple independent LLM-based agents on QA task for unknown environment. While other methods in the field of Mixture of Experts have used multiple LLM-based models for QA tasks, we specifically investigate this paradigm with a dependence on embodied exploration in a household environment. We compare the accuracy of various aggregation methods: majority voting, debating, and CAM. We show that various machine learning methods are suitable to train CAM and consistently outperform the non-learning-based aggregation methods.
A key benefit of CAM is that it requires no communication during inference, reducing time costs. We also show that CAM is more robust to malicious agents than the majority vote and debate methods. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
7 Limitations and Future Work
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions. It would also be useful to remove the need to explicitly mention the target object in the query. Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding. Third, our results hinge on the quality of the observation data. In our work, we relied on the GLIP model. Future work should analyze the effect of scene understanding algorithms on our setup.
Appendix A Code and Dataset
@@ Appendix B Multi-LLM vs LLM
Appendix C Additional Visualization of Results
@@
Figure 9: Visualization of Decision Tree from Section 5.
Appendix D Decision Tree Visualization
In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section 5. This tree in Figure 9 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents. One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations. We see that the main feature used was actually the object query. We provide the object encoding labels in Table 5.